Discuss why managers estimate a cost function and use Cost volume Profit analysis? Give numerical example of cost function and Cost Volume Profit Analysis and analyze how it will be used by managers?

           b- Suppose actual costs are higher than estimated cost. Analyze why you may have this difference between actual and estimated costs?
Answer #1

Managers estimate a cost function and use Cost volume Profit analysis due to following reasons-

  • A cost function is a function of input prices and output quantity whose value is the cost of making that output given those input prices, often applied through the use of the cost curve by companies to minimize cost and maximize production efficiency.

  • Cost function

    C(x) = F + Vx

    C = Total cost

    F = Fixed cost

    V = Variable cost per unit

    x = Number of units produced

  • Cost-volume-profit (CVP) analysis is a method of cost accounting that looks at the impact that different levels of costs and volume have on operating profit. This analysis, also commonly known as break-even analysis, looks to determine the break-even point for different sales volumes and cost structures, which can be useful for managers making short-term economic decisions.

So, managers estimate a cost function and use Cost volume Profit analysis to

firstly estimate cost through Cost function &

then use Cost volume profit analysis to

achieve the required sales to get

desired profits in the business.

Example of cost function-

'Y Co.' has two product lines. It manufactures 'a' & 'b'. While preparing its budget for the following year, it uses the cost function formula to determine the ideal product mix.

The company pays a fixed amount of $18,000 every year towards rent, insurance, and machinery related costs.

The variable costs that Mollin Industries will incur for its two product lines are:

'A': $1.75 per foot.

'B': $3 per foot.

70,000 feet of 'A' or 40,000 feet of 'B' an be produced in a year.

The cost function formula yields the following calculation:

'A'- C(x) = F + Vx=18,000+1.75(70,000 feet)=$140,500

'B'- C(x) = F + Vx=18,000+3(40,000 feet)=$138,000

cost function formula allows a manager to calculate its total cost of production in different manufacturing scenarios

Example of Cost Volume Profit Analysis-

'X' company desired an accounting profit of $90,000, where fixed cost is $70,000, contribution margin of 40%, the total sales revenue is found by dividing $160,000 (the sum of fixed costs and desired profit) by the contribution margin of 40%. This example yields a required sales revenue of $400,000.

The CVP formula can be used by managers to calculate the sales volume needed to cover costs and break even.

b-Analysis why we may have difference between actual and estimated costs-

  • Before doing any activity estimated costs is calculated through various ways like from past experience, through market quotes, estimation etc.
  • Estimated costs are not exact as this was probable costs, so when we actually incurr the expense then it may differ.
  • So, this is the reason of the difference.

Thank You!

Please give your feedback